# Multi-modal and Multi-scale Spatial Environment Understanding for Immersive Visual Text-to-Speech

### v1

@@ Introduction
To evaluate the effectiveness of selecting depth features using shared Topk indices from the RGB image and to compare this approach with independent semantic-guided methods, we focus on the efficacy of index-sharing and the impact of different Topk values. Specifically, we design experiments to compare two feature selection strategies: shared Topk indices and independent semantic-guided selection.
In the shared index strategy, Topk critical regions from RGB images guide depth feature selection. In contrast, the semantic-guided strategy independently selects features from RGB and depth images based on their respective semantic information. Various Topk values (e.g., 20, 40, …, 240) are tested to observe their effects on performance, evaluated using two objective metrics, with lower values indicating better performance.
The experimental results demonstrate that the shared Topk strategy consistently outperforms the independent semantic-guided strategy for all tested values of Topk. This approach yields lower values for objective metrics, indicating more natural audio generation that better aligns with the environment. As the Topk value increases, performance improves as more comprehensive spatial information is captured, though it plateaus or declines slightly after a threshold (e.g., 140).

### v2

@@ Introduction
To evaluate the effectiveness of selecting depth features using shared Topk indices from the RGB image and to compare this approach with independent semantic-guided methods, we focus on the efficacy of index-sharing and the impact of different Topk values. Specifically, we design experiments to compare two feature selection strategies: shared Topk indices and unshared Topk indices.
In the shared indices strategy, Topk critical regions from RGB images guide depth feature selection. In contrast, the unshared Topk indices strategy independently selects features from RGB and depth images based on their respective semantic information. Various Topk values (e.g., 20, 40, …, 240) are tested to observe their effects on performance, evaluated using two objective metrics, with lower values indicating better performance.
The experimental results demonstrate that the shared Topk strategy consistently outperforms the unshared Topk strategy for all tested values of Topk. This approach yields lower values for the objective metrics, indicating more natural audio generation that better aligns with the environment. As the Topk value increases, performance improves as more comprehensive spatial information is captured, though it plateaus or declines slightly after reaching a threshold (e.g., 140).